• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M Jones

    Log entry time 16:00:30 on November 06, 2006

    Entry number 189002

    keyword=Day shift summary

    8-10 Production running on 4He.
    10-1:20 Beam down for Mott and beam studies
    10-12:30 Controlled access ( Jack Segal, Mark Stevens, Sirish and Alex
    in hall
    12:00 Cryo problems cause the target to warm up.
    1:30 Beam back and running with 12C target.
    Target slowly recoverying at the end of shift.